OpenStudy (anonymous):

Solve for h. A = ((a+b)/2)h

OpenStudy (anonymous):

h=A-(a+b)/2

OpenStudy (anonymous):

? h is times all that not subtract how do you get that answer?

OpenStudy (anonymous):

oh! my bad, hold on

OpenStudy (anonymous):

so far i have A=[h(a+b)]/2 after multiplying the (a+b)/2 by the h

OpenStudy (anonymous):

Oh cool that does it. 2A = h(a+b) 2A/(a+b) = h cool thanks

OpenStudy (anonymous):

:) ok